Gator Beach Marts, a chain of convenience grocery stores in the Fort Lauderdale area, has store hours that fluctuate from month to month as the tourist trade in the community varies. The utility costs for one of the company’s stores are listed below for the past six months.

  

Month Total Hours
of Operation
Total Utility
Cost
January 550 $ 1,620
February 600 1,700
March 700 1,900
April 500 1,600
May 450 1,350
June 400 1,300

Required:

1-a. Use the high-low method to estimate the cost behavior for the store’s utility costs. Express the cost behavior in formula form (Y= a + bX).

Monthly utility cost =

1-b. What is the variable utility cost per hour of operation?

Variable cost per hour

3-a. Construct an Excel spreadsheet and use the Excel commands to perform a least-squares regression. Estimate the cost behavior for the store's utility cost. Express the cost behavior in formula form.

Monthly utility cost =

3-b. What is the variable utility cost per hour of operation?

Variable utility cost per hour

4. During July, the store will be open 300 hours. Predict the store’s total utility cost for July using each of the cost-estimation methods employed in requirements (1) and (3).

Cost Predictions High-low method Least-square regression
Answer #1

1-a
Variable cost per unit = ($1900-1300)/(700-400) = $2 per hour
Fixed Cost = $1900 - 700 x 2 = $500
Y = $500 + 2 X

1-b
Variable utility cost per hour = $2

3-a

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.97559
R Square 0.951776
Adjusted R Square 0.93972
Standard Error 54.90902
Observations 6
ANOVA
df SS MS F Significance F
Regression 1 238023.3 238023.3 78.94638 0.000886
Residual 4 12060 3015
Total 5 250083.3
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 501 123.3054 4.063083 0.01531 158.6494 843.3506 158.6494 843.3506
Total Hours of Operation 2.02 0.227345 8.885178 0.000886 1.388789 2.651211 1.388789 2.651211

Y = 501 + 2.02 X

3-b
Variable cost per hour = $2.02 per hour

4
(1) Total Cost = $500 + 300 x $2 = $1100

(2) Total Cost = $501 + 300 x $2.02 = $1107